OpenStudy (anonymous):

Marilyn uses a credit card with a 19.9% APR compounded monthly to pay for car repairs totaling $991.38. She can pay $410 per month on the card. What will the total cost of this purchase be? a) $1,021.01 b) $1,188.66 c) $991.38 d) $1,192.39

OpenStudy (anonymous):

It didn't really say anything about how to do this kind of problem, though =/ thank you, nontheless

OpenStudy (mathstudent55):

$991.38 is charged to her card on day 1 After 1 month, she owes $991.50 + one month's interest on $991.50 One month's interest is: 19.9/100/12 x $991.5 = $16.44 So after the first month she owes $991.50 + $16.44 = $1007.94 She pays $410, so her balance now is $1007.94 - $410 = $597.94 After one more month, her new balance is $597.94 + one month's interest on $597.94 One month's interest is: 19.9/100/12 x $597.94 = $9.92 So after the 2nd month she owes $597.44 + $9.92 = $607.86 She pays $410, so her balance now is $607.86 - $410 = $197.86 After one more month, her new balance is $197.86 + one month's interest on $197.86 One month's interest is: 19.9/100/12 x $197.86 = $3.28 So after the 2nd month she owes $197.86 + $3.28 = $201.14 She pays $201.14, and pays it off. The total she paid was $410 + $410 + $201.14 = 1021.14

OpenStudy (anonymous):

thanks a bunch!

jimthompson5910 (jim_thompson5910):

Month 1: 991.38*(1 + 0.199/12) = 1007.82 Make payment: 1007.82 - 410 = 597.82 -------------------------------------------------------------- Month 2: 597.82*(1 + 0.199/12) = 607.73 Make payment: 607.73 - 410 = 197.73 -------------------------------------------------------------- Month 3: 197.73*(1 + 0.199/12) = 201.01 Make payment: 201.01 Notice how the balance is less than the monthly payment. So this is when the card will be fully paid off. -------------------------------------------------------------- Payments made: 410, 410, 201.01 Add up the payments to get 410+410+201.01 = 1,021.01

jimthompson5910 (jim_thompson5910):

So the total cost is $1,021.01

OpenStudy (anonymous):

thank you--much appreciated!

jimthompson5910 (jim_thompson5910):

yw

OpenStudy (mathstudent55):

I just noticed that I used $991.50 in my fisrt calculation even though I had written $991.38 originally. I just redid it with the correct $991.38. $991.38 is charged to her card on day 1 After 1 month, she owes $991.38 + one month's interest on $991.38 One month's interest is: 19.9/100/12 x $991.38 = $16.44 So after the first month she owes $991.38 + $16.44 = $1007.82 She pays $410, so her balance now is $1007.82 - $410 = $597.82 After one more month, her new balance is $597.82 + one month's interest on $597.82 One month's interest is: 19.9/100/12 x $597.82 = $9.91 So after the 2nd month she owes $597.82 + $9.91 = $607.73 She pays $410, so her balance now is $607.73 - $410 = $197.73 After one more month, her new balance is $197.73 + one month's interest on $197.73 One month's interest is: 19.9/100/12 x $197.73 = $3.28 So after the 2nd month she owes $197.73 + $3.28 = $201.01 She pays $201.01, and pays it off. The total she paid was $410 + $410 + $201.01 = 1021.01